Helpdesk Tickets. There have been 16 helpdesk tickets submitted thus far (mostly straight forward questions). There have been 50+ udb-admin tickets (mostly institution/password requests). The PST server load has been rather low. If RickLively can find a good time he may kick the server.
PST Review Items. There is a total of six PST review tickets for 17A. Two have been released for testing and four are in progress. Of the two released one has been tested and is done. DanaBalser will check the cookbook to make sure there are no issues with starting the review process next week.
PHT Requirements. RickLively will update the status of 16B PHT tickets and then GustaafVanMoorsel will inspect this list. BarryClark has made the following changes to the VLA Prioritizer: converted about 15 internal parameters to a properties file; added the ability to control the amount of high frequency time that is assigned to filler; added the ability to control the amount of filler assigned for projects that have been assigned some A (or B) time. PaulMarganian was working on importing GBT proposals with MUSTANG2.0 or Vegas Pulsar resources from the PHT to the GB PHT. He initially had problems but ToneyMinter thinks this has been fixed.
SRP Recruitment. MarkClaussen has secured the ETP SRP vacancy. No word from JayLockman on the remaining EGS SRP vacancy. It is not clear if we have a 17A chair for the TAC.